Find the solution set of the given equation over the set of complex numbers. (complete the square).
x^2=10x-34

Answer by checkley77(12844)

x^2=10x-34
X^2-10X=-34 COMPLETE THE SQUARE.
(X^2-10X+25)=-34+25
(X+5)^2=9 TAKE THE SQUARE ROOT OF BOTH SIDES.
X+5=3
X=3-5
X=-2 ANS.